OpenStudy (anonymous):

logx+log3=log18 x=15 is this right?

imqwerty (imqwerty):

no that not correct what did u do?

OpenStudy (anonymous):

take the log out and find X........just like the teacher said

OpenStudy (anonymous):

x+3=18 sub 3 from both side and end up with x=15

imqwerty (imqwerty):

Lol yes but that doesn't means that u will remove the log terms okay we will do it step wise :) 1st try to use this rule on the left side-> \(\large \color{blueviolet}{log(a)+log(b)=log(a \times b)}\)

OpenStudy (anonymous):

ok?

OpenStudy (triciaal):

@williettelover @imqwerty response in words when you add logs you multiply the numbers

OpenStudy (anonymous):

it just says to solve for x im confused

OpenStudy (triciaal):

this is still an equation the left side is equal to the right side

OpenStudy (triciaal):